# SATURN: SAT-based Reinforcement Learning to Unleash Language Model Reasoning

### v1

@@ 3.6 RQ4: SATURN Enhances Self-verification in LLMs’ Reasoning Trajectories
RQ4 investigates whether SATURN affects LLMs’ reasoning behavior. On Qwen2.5-7B-Instruct-1M, we observe a gradual increase in response length during training, as illustrated in Figure 4, replicating the lengthening phenomenon reported in the R1 and Logic-RL [12, 45].
@@
To examine whether such reasoning patterns generalize, we present case studies across SAT and math domains. Figure 12 shows that solving SAT variables requires rechecking all clauses, naturally encouraging self-verification. In Figure 5, SATURN-7B verifies intermediate conclusions within a small scenario and successfully chooses the correct solution path. In contrast, the baseline model reaches a wrong answer and skips verification, even when inconsistencies are detected. More reasoning trajectories are provided in Appendix K. These results suggest that the self-verification patterns learned from SAT transfer well to math and programming tasks, improving reasoning robustness and reliability.
@@ 4.1 Limitations of Reasoning Capability Learned from SATURN
During curriculum learning, we observed that as the number of training iterations increases, the improvements in math and programming tasks tend to plateau, which is consistent with the findings in Logic-RL. Detailed evaluation results are provided in Appendix I. This plateau may stem from several factors: ❶ Knowledge limitations. SATURN improves formal logical reasoning but does not provide domain-specific knowledge supervision. This limits its effectiveness in tasks requiring mathematical or algorithmic knowledge. ❷ Context window bottlenecks. SAT problems are NPC tasks, and the required reasoning length grows exponentially with increasing problem difficulty. This leads to bottlenecks in the model’s capability to handle increasingly complex tasks. ❸ Limited plasticity and forgetting. Model plasticity and catastrophic forgetting are known limitations that hinder further improvements with additional training stages [5, 14].
@@ 4.2 Potential of SATURN on Stronger Models
To explore the potential of SATURN to stronger models, we evaluate frontier LLMs on SAT tasks using the extended SAT instances. Results are shown in Appendix J. Although these LLMs exhibit stronger performance, they still make common errors such as hallucinating clauses, confidently committing to incorrect decisions, or failing to apply basic logical rules. Even more advanced LLMs still struggle to solve complex SAT problems. We believe SATURN remains a promising approach for enhancing reasoning in stronger LLMs. With sufficient computation, SATURN can offer a scalable, verifiable, and controllable path to further improve reasoning capabilities.
@@ 5.1 Constructing Reasoning Tasks for Reinforcement Learning
Several works have explored constructing reasoning tasks to improve the reasoning capability of LLMs. Logic-RL [45] and LMRL Gym [6] train LLMs on natural language logic puzzles but lack scalability due to their limited puzzle set. ScaleQuest [13], Entity-Deducing Game [50], and K&K [44] propose automatic generation of constructing questions, but rely on LLM sampling or handcrafted templates, making large-scale generation costly and hard to verify. CodeDPO [49] and PuzzBench [39] employ LLM-based verification, which may fail silently and cannot ensure correctness. Wolf Game [43, 46] focus on multi-step logic reasoning but offer no control over task difficulty, limiting their support for curriculum learning. Overall, these tasks fall short in scalability, verifiability, or controllable difficulty. See Appendix L for detailed comparisons.
@@ 5.2 SAT-Based Evaluation of LLM Reasoning Capability
Recent studies have evaluated the reasoning capability of LLMs on SAT problems. Most of these works focus on analyzing model behavior around the SAT phase transition [18, 32, 29], where problem hardness peaks. However, the phase transition theory is originally designed for heuristic SAT solvers and does not align well with the reflective and verification-based reasoning processes of humans or LLMs. These studies also lack a fine-grained scalable difficulty framework and typically divide difficulty based on the phase transition threshold. They are further limited to supervised fine-tuning and do not consider large reasoning models with long-CoT reasoning capability trained via RL. Our work addresses these limitations by building a progressive evaluation and curriculum learning pipeline, enabling precise difficulty control and the generalization of LLMs.

### v3

@@ 3.6 RQ4: SATURN Enhances Self-verification in LLMs’ Reasoning Trajectories
RQ4 investigates whether SATURN affects LLMs’ reasoning behavior. On Qwen2.5-7B-Instruct-1M, we observe a gradual increase in response length during training, as illustrated in Figure 4, replicating the lengthening phenomenon reported in the R1 and Logic-RL [deepseek_r1, Logic-RL].
@@
To examine whether such reasoning patterns generalize, we present case studies across SAT and math domains. Figure 12 shows that solving SAT variables requires rechecking all clauses, naturally encouraging self-verification. In Figure 5, SATURN-7B verifies intermediate conclusions within a small scenario and successfully chooses the correct solution path. In contrast, the baseline model reaches a wrong answer and skips verification, even when inconsistencies are detected.
Recent studies [Cognitive, Meta-Abilities] identify core behaviors shared by expert human reasoners and LLMs, such as verification and backtracking. These behaviors are domain-agnostic and provide fundamental reasoning patterns applicable to a wide range of tasks. In line with these findings, SATURN reinforces similar behaviors during SAT solving, leading to more structured reasoning trajectories. More reasoning trajectories are provided in Appendix J to illustrate how SATURN works. These results suggest that the self-verification patterns learned from SAT transfer well to math and programming tasks, improving reasoning robustness and reliability.
@@ 4.1 Limitations of Reasoning Capability Learned from SATURN
During curriculum learning, we observed that as the number of training iterations increases, the improvements in math and programming tasks tend to plateau, which is consistent with the findings in Logic-RL. Detailed evaluation results are provided in Table 5. This plateau may stem from several factors: ❶ Knowledge limitations. SATURN improves formal logical reasoning but does not provide domain-specific knowledge supervision. This limits its effectiveness in tasks requiring mathematical or algorithmic knowledge. ❷ Context window bottlenecks. SAT problems are NP-complete tasks, and the required reasoning length grows exponentially with increasing problem difficulty. This leads to bottlenecks in the model’s capability to handle increasingly complex tasks. ❸ Limited plasticity and forgetting. Model plasticity and catastrophic forgetting are known limitations that hinder further improvements with additional training stages [forgetting1, forgetting2].
Model AIME 24/25 AMC 22/23 Math500 GPQA-D LiveCodeBench Avg.
DeepSeek-R1-Distill-Qwen-1.5B 21.6 65.1 83.6 30.3 16.4 43.4
SATURN-1.5B-Iteration-1 26.7 68.6 85.0 33.3 16.9 46.1
SATURN-1.5B-Iteration-2 28.3 73.5 84.6 37.4 17.4 48.2
SATURN-1.5B-Iteration-3 28.3 66.3 85.8 36.9 16.7 46.9
DeepSeek-R1-Distill-7B 50.0 80.7 93.2 49.0 35.4 61.7
SATURN-7B-Iteration-1 48.3 83.1 94.6 50.5 36.6 62.6
SATURN-7B-Iteration-2 48.3 85.5 95.0 50.5 37.7 63.4
SATURN-7B-Iteration-3 46.7 87.9 93.2 58.1 38.1 64.8
@@ 4.2 Potential of SATURN on Stronger Models
To explore the potential of SATURN to stronger models, we evaluate frontier LLMs on SAT tasks using the extended SAT instances. Results are shown in Appendix I. Although these LLMs exhibit stronger performance, they still make common errors such as hallucinating clauses, confidently committing to incorrect decisions, or failing to apply basic logical rules. Even more advanced LLMs still struggle to solve complex SAT problems. We believe SATURN remains a promising approach for enhancing reasoning in stronger LLMs. With sufficient computation, SATURN can offer a scalable, verifiable, and controllable path to further improve reasoning capabilities.
@@ 5.1 Constructing Reasoning Tasks for Reinforcement Learning
Several works have explored constructing reasoning tasks to improve the reasoning capability of LLMs. Logic-RL [Logic-RL] and LMRL Gym [lmrl] train LLMs on natural language logic puzzles but lack scalability due to their limited puzzle set. ScaleQuest [question_synthesis], Entity-Deducing Game [20_Question], and K&K [kk2] propose automatic generation of constructing questions, but rely on LLM sampling or handcrafted templates, making large-scale generation costly and hard to verify. CodeDPO [codedpo] and PuzzBench [puzzbench] employ LLM-based verification, which may fail silently and cannot ensure correctness. Wolf Game [wolf1, wolf2] focus on multi-step logic reasoning but offer no control over task difficulty, limiting their support for curriculum learning. Overall, these tasks fall short in scalability, verifiability, or controllable difficulty. See Appendix K for detailed comparisons.
@@ 5.2 SAT-Based Evaluation of LLM Reasoning Capability
Recent studies have evaluated the reasoning capability of LLMs on SAT problems. Most of these works focus on analyzing model behavior around the SAT phase transition [sat_iclr_1, sat_iclr_solver, sat_blog], where problem hardness peaks. However, the phase transition theory is originally designed for heuristic SAT solvers and does not align well with the reflective and verification-based reasoning processes of humans or LLMs. These studies also lack a fine-grained scalable difficulty framework and typically divide difficulty based on the phase transition threshold. They are further limited to supervised fine-tuning and do not consider large reasoning models with long-CoT reasoning capability trained via RL. Our work addresses these limitations by building a progressive evaluation and curriculum learning pipeline, enabling precise difficulty control and the generalization of LLMs.
@@ 6 Conclusion and Future Work
Acknowledgements
This research is supported by the National Key R&D Program under Grant No. 2023YFB4503801, the National Natural Science Foundation of China under Grant No. 62192733, 62192730, 62192731, the Major Program (JD) of Hubei Province (No.2023BAA024).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the main contributions, including the curriculum learning framework SATURN and the SATURN-2.6k dataset. All supported by experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of SATURN in Section 4.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Any informal proof provided in the core of the paper is complemented by formal proofs in the appendix, including the rigorous derivation of SAT estimation.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We release all code, models, datasets, and experimental results to ensure full reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We release all code, models, datasets, and experimental results, along with detailed information on the execution environment to ensure full reproducibility.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide all important parameters needed to understand the results in the appendix, and the code scripts include complete training and test details.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Following current evaluation practices in recent large model benchmarks, we report the unbiased estimator of the pass@k metric. This expectation-based metric is designed to reflect statistical significance across multiple samples and inherently captures sampling variability. Although we do not report explicit error bars, pass@k inherently reflects statistical reliability. Further details are provided in Appendix E.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide details on the GPU resources used for experiments as well as information on all data storage.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed and ensured that all research conducted in this paper fully conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper focuses on improving LLMs’ reasoning capability through SATURN curriculum learning framework, without direct deployment or application, and thus does not raise societal impacts concerns.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no foreseeable risks of misuse, as it does not involve the release of high-risk models or datasets.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly credit all used assets in the paper and appendix, including code, datasets, and models. Licenses and terms of use are explicitly stated and respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release new code, datasets, and models. All new assets are thoroughly documented, including usage instructions and data format specifications, which are provided alongside the release.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve any form of crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing or research with human subjects, so IRB approval is not applicable.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This paper does not involve the use of LLMs as an important, original, or non-standard component of the core methodology.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.